# StructSum Generation for Faster Text Comprehension

### v1

@@ 1 Introduction
Figure 1: Overview of (a) tables and (b) mind map generation prompts. Prompting steps are colored. Figure (a) show divide-and-generate prompt for multiple table generation. Input passages is first segmented into sub-passages, followed by table generation. Figure (b) is iterative expansion based mind map prompt. Starting from main concept generation, at each iteration current mind map is expanded until termination.
@@
  • Tables are well-studied in the NLP literature. However the vast majority of the work focused on simpler tasks where tables are inputs – such as QA (Herzig et al., 2020), semantic parsing (Bogin et al., 2019), NLG Andrejczuk et al. (2022); Puduppully and Lapata (2021); Laha et al. (2020), etc. – rather than outputs. Indeed, faithfully transforming an arbitrary text passage into a table is a difficult task as the model must deal with different challenges, such as reasoning at multiple levels, dealing with missing information, and visually consistent formatting. Motivated by the limitations above, we propose to generate multiple tables instead. We argue that this is a simpler task for an LLM, as shown in Figure 2, which compares single-table and multi-table generation side by side. We therefore propose a divide-and-generate prompting approach that first divides the input text into multiple text passages, each representing a sub-topic, followed by an LLM prompt to generate a table-caption pair for each smaller passage. This decomposition allows the model to generate smaller, focused and more informative tables, especially for complex text passages with multiple sub-topics.
@@
  • Mind maps (Hu et al., 2021; Wei et al., 2019) are less studied in the literature, but are helpful for comprehension and learning (Buzan, 1996; Dhindsa et al., 2011). Mind maps are complementary to tables in their structure, allowing for more flexibility and dynamism than tables, as they are inherently schema-less. However, generating mind maps with LLMs presents several challenges: (i) the model first need to select a central concept, that is the fulcrum of all the successive extractions, as mind maps revolve around a central root node; (ii) being a schema-less abstraction, each connecting branch has its own independent sub-topic, making it difficult to automatically add branches all at once; (iii) to ensure readability and well-structuredness each leaf node should terminates the path in a way that concludes the idea or sub-topic; (iv) depending on the information density, some paths may be shorter than others. Therefore, the model should decide whether or not a branch is worth expanding. Following the structure of these observations, we propose an iterative prompting technique for mind map generation. We initialize the mind map by generating the root concept. At each iteration, we decide either to expand the current mind map further or stop the process. During the expansion step, we prompt the model to add branches to the current leaf nodes. We represent the mind map as a JSON object, as it is easy to parse and verify.
@@
Through extensive experimentation with PaLM-2 (PaLM2, 2023), we show that LLMs are not always effective at generating StructSums that are factual and structurally correct. To overcome these issues we propose a pipeline for structured data generation. Our pipeline consists of structure-specific prompts followed by critics to assess output quality along three different dimensions, that are common both to tables and mind maps: (i) Factuality, (ii) Local Structure and (iii) Global Structure. We found that our proposed critics improved the overall quality of the generated output by $37$pp for mind maps and $15$pp for tables.
@@ 2 Related Work
Due to the cost of human evaluation, LLMs are used to critic the generated outputs (Wang et al., 2023). Recent instructions tuned models, such as, GPT-4 (OpenAI, 2023) and ChatGPT (OpenAI, 2022) are shown to be strong evaluators. To avoid using external APIs,  Kim et al. (2023); Wang et al. (2023) fine-tune a smaller pretrained model to critic model responses. We are interested in evaluating the quality structured outputs using critics and self-correct based on the feedback. As a part of data generation pipeline, our focus is on filtering instances that are incomplete and are not factually grounded.
@@ 3.1 Tables: Divide & Generate prompting
To address these limitations, we propose a divide-and-generate approach that dynamically partitions the passage into smaller subtopic segments. While deterministic rule-based chunking methods (e.g., based on word or sentence count) can be employed, they often produce suboptimal results due to potential under-chunking, over-chunking, and the absence of division for certain instances. Therefore, the chunking must be adaptive and depend on the input text and its sub-topic distribution. We use a one-shot prompt for this step, as shown in Appendix C (Figure  14). After the chunking, we prompt the model to generate a table along with its caption for each sub-passage obtained in the previous step.
@@ 3.2 Mind maps: Iterative Prompting
Contrary to tables, mind maps are more flexible and present a different set of challenges. The first challenge is representation. We desire a representation that is (i) close to a familiar format, and (ii) is easily parsable and verifiable using current tools. JSON meets both of these requirements. The second challenge is that mind maps, unlike tables where each row can be produced linearly, necessitate attaching information in different locations depending on which branch is being expanded. This requires the model to think radially.
We propose an iterative prompting for mind maps generation. Algorithm 1 shows the overall procedure. Details of each prompt is in Appendix C. We start by generating the root concept that becomes the central node for the mind map. This separate step allows the model to independently reason about the theme of the passage. After generating the root, at each step we prompt the model to decide if current mind map can be expanded further. If the model decides to expand (line 5), we prompt the model using the current mind map to add more branches. Otherwise, the procedure terminates and we return the current mind map. At each expansion step we sample multiple mind maps. Utilizing the fact that JSON verification is cheaper we select the topmost JSON that is parsed correctly. In the rare case, when none of the samples are parsable we call a critic prompt to correct the top JSON (line 7).
@@ 4.3 Global Structure Critic
For tables, we simply verify whether the table is well formatted: e.g. we verify equal number columns in the header and subsequent rows, therefore ignoring semantic content of the table and only focusing on form rather than the content. This is realized via simple heuristics.
@@ 5 Semantic Coverage using Auto-QA
In this section we propose an automatic way to assess the quality and the general usefulness of StructSums introducing Auto-QA coverage as proxy metric. This metric measures the semantic coverage or percentage of questions that are answerable when using a StructSum   $s$, instead of the full text passage $t$. Formally it is defined as:
@@
where $\text{GenQA}(x)$ is a function that generates $(q,a)$ pairs given the input text passage $t$, $Q(s,q_{i})$ is a function that generates an answer given in input a StructSum $s$ and the question $q_{i}$, whereas the indicator function $\mathbbm{1}_{E_{a_{i}}}(x)$ asses the answer equivalence between $a_{i}$ and $x$. Figure 18 in Appendix C, show all the prompts associated with AutoQA module (Deutsch et al., 2021; Fabbri et al., 2022).
@@ 7 Dataset
To test our pipeline on a diverse set of input passages, we selected Wikipedia text as the source. Specifically, we started with the English section of Wiki40b (Guo et al., 2020) dataset. The dataset is cleaned up by page filtering to remove disambiguation pages, redirect pages, deleted pages, and non-entity pages. We then iterated over the dataset to extract passages that are inputs to our prompt.
@@ 8.1 Quality impact of prompting style and automated critics
Table 2: Table / mind map accuracy per prompt style. Outputting multiple tables provides higher quality for the table modality. For mindmaps, an iterative approach is to be preferred to a CoT approach. Full prompts are reported in the Appendix C.
@@
Table 2 show the results for both the modalities. For table generation task we find that annotators prefer multiple tables generation outputs compared to single table generation. For mind map, we compare chain-of-thought (Wei et al., 2022) with our proposed iterative generation strategy described in Algorithm 1. We find that iterative generation were preferred over simpler prompt outputs.
@@ 8.2 Do Critics Align with Human Ratings?
Through our human annotations results in Section 8.1, we find that many generated outputs are not of acceptable quality. To improve the quality of the generated data and to avoid costly human annotations, we propose to use a combination of critics as a measure of data quality. To verify the efficacy of our critics, we first filtered the generated dataset with our critics. Specifically, we performed a logical AND of individual critics and filtered the instances that do not pass the criterion. We then conducted the same evaluation of Section 8.1.
@@ 8.4 Is Auto-QA a reasonable metric?
Table 4 shows the overall results. Second (Mind maps) and third (Multi-Table) row show the comparison between Human QA and Auto QA for our proposed divide-and-generate prompt for tables and iterative prompt for Mindmap generation. We can see that Auto-QA has comparable results and is a reasonable substitute for human generated questions as a measure of semantic coverage.
@@ 8.5 Multiple Tables vs Single Table
To check whether generating multiple tables is better at covering more information, we perform a comparison between the ability to answer question by generating single or multiple tables. On comparing Multi-Table and Single Table row in Table 4, we observe that for both Auto-QA and Human QA generating multiple table provides more coverage. So in addition to the benefits such as comparatively better verifiability and robust generation, multiple table generation are also better at covering more semantic information.
@@ 8.6 Query Focused Generation
In many cases user intent is known in advance, for example, a user query to search or LLM-based Assistant interface (e.g., ChatGPT, Bard, etc.). We explore the possibility of generating structured data in the presence of a query. We perform a preliminary analysis by adding the query in single table generation prompt. As we can see in last row in Table 4, query focused generation improve the performance by more than $20$ for Auto-QA and $25$ points for Human-QA. Since this need further investigation in terms of prompting and more importantly detailed analysis of output quality. We leave a comprehensive exploration for query focused structured data generation as future work.
@@ 8.7 Are StructSums useful?
We evaluate whether StructSum are useful abstractions for the users. For this we design a timed text comprehension based user study. We assume that user is looking to answer a specific query, i.e. has a specific intent. We measure time taken to satisfy the user intent as a proxy of usefulness.
@@
StructSums for the study were generated using our data generation pipeline and critic-based filtering, as discussed in Section 4. In total $600$ instances were annotated, equally divided into different context combinations for mind maps and table generation. Annotators consistently answered correctly across all context combinations (Appendix B), suggesting that the level of context did not significantly impact their accuracy.
Figure 5 shows the overall results. The plots show 95% confidence interval of time taken by the annotators when using different modality:
@@
We note that $\langle q,s+t\rangle$ performs worse than $\langle q,s\rangle$, we believe this is due to the fact that the annotators cross-checked the answer from both the modalities. This led to increase in time taken to answer the question.
9 Conclusion
In this work we studied the potential of structured representations like tables and mind maps to enhance information comprehension. Utilizing our divide-and-generate prompting and iterative expansion, we achieved significant improvements in output quality ($+37pp$ for mind maps, $+15pp$ for tables) using structure-specific prompts and critics. We proposed Auto-QA based coverage metric that automatically generates QA pairs from the input text and uses the StructSum outputs to answer them.
10 Acknowledgment
We are grateful to Massimo Nicosia and Yasemin Altun for their feedback and comments on the paper.
Limitations
We outline the limitations of our work to ensure transparency and inspire future research. First, the structured output representations we experimented with are limited to tables and mind maps. However, to comprehensively evaluate the effectiveness of our critics and pipeline, it is desirable to also evaluate other input and output modalities, e.g. image and video, considering the recent advances in VLMs. Secondly, our work and experimental findings are limited to only English sources. We plan to also explore multilingual structured summaries in future works. Third, we would to warn against the risk of blindly trusting models to generate structured summaries from an input accurately. Although we take extra care to increase the factuality of the outputs via the use of critics, and experimentally validate QA coverage, we believe that special care should be taken to verify outputs in accuracy-sensitive applications. Despite these limitations, our work serves as an initial step in constructing reliable structured summarization evaluations, models and applications. We hope future research can greatly benefit from this starting point.
Appendix A Data Generation statistics
@@
(a) Input text for table genetation.
(b) Single table generation output.
(c) Multiple table generation output.
Appendix B User Study

### v2

@@ Abstract
StructSum Generation for Faster Text Comprehension
Parag Jain1thanks:   Work done while interning at Google DeepMind.  , Andreea Marzoca2, Francesco Piccinno2 School of Informatics, University of Edinburgh1 Google DeepMind2 parag.jain@ed.ac.uk, {andreeam,piccinno}@google.com
@@ 1 Introduction
Figure 1: Overview of (a) tables and (b) mind map generation prompts. The prompting steps are colored. Figure (a) illustrates the divide-and-generate prompt. The input passage is initially segmented into sub-passages, followed by the generation of multiple tables. Figure (b) demonstrates the generation process for mind maps. After the main concept has been generated, an iterative expansion phase ensues, during which the mind map is expanded until termination.
@@
  • Tables are well-studied in the NLP literature. However the vast majority of the work focused on simpler tasks where tables are inputs – such as QA (Herzig et al., 2020), semantic parsing (Bogin et al., 2019), NLG Andrejczuk et al. (2022); Puduppully and Lapata (2021); Laha et al. (2020), etc. – rather than outputs. Indeed, faithfully transforming an arbitrary text passage into a table is a difficult task as the model must deal with different challenges, such as reasoning at multiple levels, dealing with missing information, and visually consistent formatting. Motivated by the limitations above, we propose to generate multiple tables instead. We argue that this is a simpler task for an LLM, as shown in Figure 2, which compares single-table and multi-table generation side by side. We therefore propose a divide-and-generate prompting approach (see Figure 1) that first divides the input text into multiple text passages, each representing a sub-topic, followed by an LLM prompt to generate a table-caption pair for each smaller passage. This decomposition allows the model to generate smaller, focused and more informative tables, especially for complex text passages with multiple sub-topics.
@@
  • Mind maps (Hu et al., 2021; Wei et al., 2019) are less studied in the literature, but are helpful for comprehension and learning (Buzan, 1996; Dhindsa et al., 2011). Mind maps are complementary to tables in their structure, allowing for more flexibility and dynamism than tables, as they are inherently schema-less. However, generating mind maps with LLMs presents several challenges: (i) the model first need to select a central concept, that is the fulcrum of all the successive extractions, as mind maps revolve around a central root node; (ii) being a schema-less abstraction, each connecting branch has its own independent sub-topic, making it difficult to automatically add branches all at once; (iii) to ensure readability and well-structuredness each leaf node should terminate the path in a way that concludes the idea or sub-topic; (iv) depending on the information density, some paths may be shorter than others. Therefore, the model should decide whether or not a branch is worth expanding. Following the structure of these observations, we propose an iterative prompting technique for mind map generation. As show in Figure 1, we initialize the mind map by generating the root concept. At each iteration, we decide either to expand the current mind map further or stop the process. During the expansion step, we prompt the model to add branches to the current leaf nodes. We represent the mind map as a JSON object, as it is easy to parse and verify.
@@
Through extensive experimentation with PaLM-2 (PaLM2, 2023), we show that LLMs are not always effective at generating StructSums that are factual and structurally correct. To overcome these issues we propose a pipeline for structured data generation. Our pipeline consists of structure-specific prompts followed by critics to assess output quality along three different dimensions, that are common both to tables and mind maps: (i) Factuality, (ii) Local Structure and (iii) Global Structure. We found that our proposed critics improved the overall quality of the generated output by +$37$pp for mind maps and +$15$pp for tables.
@@ 2 Related Work
Due to the cost of human evaluation, LLMs are used to critique the generated outputs (Wang et al., 2023). Recent instructions tuned models, such as, GPT-4 (OpenAI, 2023) and ChatGPT (OpenAI, 2022) are shown to be strong evaluators. To avoid using external APIs,  Kim et al. (2023); Wang et al. (2023) fine-tune a smaller pretrained model to critic model responses. We are interested in evaluating the quality structured outputs using critics and self-correct based on the feedback. As a part of data generation pipeline, our focus is on filtering instances that are incomplete and are not factually grounded.
@@ 3.1 Tables: Divide & Generate prompting
To address these limitations, we propose a divide-and-generate approach that dynamically partitions the passage into smaller subtopic segments. While deterministic rule-based chunking methods (e.g., based on word or sentence count) can be employed, they often produce suboptimal results due to potential under-chunking, over-chunking, and the absence of division for certain instances. Therefore, the chunking must be adaptive and depend on the input text and its sub-topic distribution. We use a one-shot prompt for this step, as shown in Appendix B (Figure  14). After the chunking, we prompt the model to generate a table along with its caption for each sub-passage obtained in the previous step.
@@ 3.2 Mind maps: Iterative Prompting
Contrary to tables, mind maps (see example in Figure 3) are more flexible and present a different set of challenges. The first challenge is representation. We desire a representation that is (i) close to a familiar format, and (ii) is easily parsable and verifiable using current tools. JSON meets both of these requirements. The second challenge is that mind maps, unlike tables where each row can be produced linearly, necessitate attaching information in different locations depending on which branch is being expanded. This requires the model to think radially.
We propose an iterative prompting for mind maps generation. Algorithm 1 shows the overall procedure. Details of each prompt is in Appendix B. We start by generating the root concept that becomes the central node for the mind map. This separate step allows the model to independently reason about the theme of the passage. After generating the root, at each step we prompt the model to decide if current mind map can be expanded further. If the model decides to expand (line 5), we prompt the model using the current mind map to add more branches. Otherwise, the procedure terminates and we return the current mind map. At each expansion step we sample multiple mind maps. Utilizing the fact that JSON verification is cheaper we select the topmost JSON that is parsed correctly. In the rare case, when none of the samples are parsable we call a critic prompt to correct the top JSON (line 7).
@@ 4.3 Global Structure Critic
For tables, we simply verify whether the table is well formatted: e.g. we verify equal number columns in the header and subsequent rows, therefore ignoring semantic content of the table and only focusing on form rather than the content. This is realized via simple heuristics implemented in Python (we do not prompt the model for these).
@@ 5 Semantic Coverage using Auto-QA
In this section we propose an automatic way to assess the quality and the general usefulness of StructSums introducing Auto-QA coverage as proxy metric.  [^1] We do not present Auto-QA as a substitute for human evaluations of quality. Instead, we propose AutoQA as a coverage metric that allows us to use synthetic question generation. This metric measures the semantic coverage or percentage of questions that are answerable when using a StructSum   $s$, instead of the full text passage $t$. Formally it is defined as:
@@
where $\text{GenQA}(x)$ is a function that generates $(q,a)$ pairs given the input text passage $t$, $Q(s,q_{i})$ is a function that generates an answer given in input a StructSum $s$ and the question $q_{i}$, whereas the indicator function $\mathbbm{1}_{E_{a_{i}}}(x)$ asses the answer equivalence between $a_{i}$ and $x$. Figure 18 in Appendix B, show all the prompts associated with AutoQA module (Deutsch et al., 2021; Fabbri et al., 2022).
@@ 7 Dataset
To test our pipeline on a diverse set of input passages, we selected Wikipedia text as the source. Specifically, we started with the English split of the Wiki40b (Guo et al., 2020) dataset  [^3] We used the version that is available via the Tensorflow datasets https://www.tensorflow.org/datasets/catalog/wiki40b.. The dataset is cleaned up by page filtering to remove disambiguation pages, redirect pages, deleted pages, and non-entity pages. Input to our prompts are passages that are obtained by splitting the Wikipedia text using the _START_PARAGRAPH_ symbol that is already provided as part of the dataset.
@@ 8.1 Quality impact of prompting style and automated critics
Table 2: Table / mind map accuracy per prompt style. Outputting multiple tables provides higher quality for the table modality. For mindmaps, an iterative approach is to be preferred to a CoT approach. Full prompts are reported in the Appendix B.
@@
Table 2 show the results for both the modalities. For table generation task we find that annotators prefer multiple tables generation outputs compared to single table generation. This can be attributed to the fact that multi-table generation enables the model to generate more concise, focused and informative tables. For mind map, we compare chain-of-thought (Wei et al., 2022) with our proposed iterative generation strategy described in Algorithm 1. We find that iterative generation were preferred over simpler prompt outputs.
@@ 8.2 Do Critics Align with Human Ratings?
Through our human annotations results in Section 8.1, we find that many generated outputs are not of acceptable quality. To improve the quality of the generated data and to avoid costly human annotations, we propose to use a combination of critics as a measure of data quality. To verify the efficacy of our critics, we first filtered the generated dataset with our critics. Specifically, we performed a logical AND of individual critics and filtered the instances that do not pass the criterion. We then sampled 100 instances from filtered examples and conducted the same evaluation as in Section 8.1.
@@ 8.4 Is Auto-QA a reasonable metric?
Table 4 shows the overall results. Second (Mind maps) and third (Multi-Table) row show the comparison between Human QA and Auto QA for our proposed divide-and-generate prompt for tables and iterative prompt for Mindmap generation. We can see that Auto-QA has comparable results and is a reasonable substitute for human generated questions as a measure of semantic coverage. We further study the limitations of Auto-QA, and the difference in Human vs Auto QA scores in Section 10.
@@ 8.5 Multiple Tables vs Single Table
To check whether generating multiple tables is better at covering more information, we perform a comparison between the ability to answer questions by generating single or multiple tables. On comparing Multi-Table and Single Table row in Table 4, we observe that for both Auto-QA and Human QA generating multiple table provides more coverage. So in addition to the benefits such as comparatively better verifiability and robust generation, multiple table generation are also better at covering more semantic information.
@@ 8.6 Query Focused Generation
In many cases user intent is known in advance, for example, a user query to search or LLM-based Assistant interface (e.g., ChatGPT, Gemini, etc.). We explore the possibility of generating structured data in the presence of a query. We perform a preliminary analysis by adding the query in single table generation prompt. As we can see in last row in Table 4, query focused generation improve the performance by more than $20$ for Auto-QA and $25$ points for Human-QA. Since this requires further investigation in terms of prompting and output quality analysis, we leave a comprehensive exploration of query-focused structured data generation as future work.
QA Type
Auto Human
Mind map 60.2 [57.0, 63.3] 61.3 [58.1, 64.4]
Multi-Table (Divide-and-generate) 72.3 [69.4, 75.2] 68.9 [66.0, 71.9]
Single Table 61.8 [58.7, 64.9] 58.1 [54.9, 61.3]
Query Focused (Single Table) 87.7 [85.5, 89.8] 85.5 [83.2, 87.8]
Table 5: Human and Auto-QA accuracy with 95% confidence interval on different modalities as context. Unlike the results presented in Table 4, we have excluded SQuAD passages for which none of the questions generated by Auto-QA passed the filter.
@@ 8.7 Are StructSums useful?
We evaluate whether StructSum are useful abstractions for the users. For this we design a timed text comprehension based user study. We assume that user is looking to answer a specific query, i.e. has a specific intent. We measure time taken to satisfy the user intent as a proxy of usefulness. Our evaluation team consists of 12 volunteers who are affiliated with our institution. Five of these volunteers are female, and seven are male. All volunteers are proficient in English, although not native speakers.
@@
StructSums for the study were generated using our data generation pipeline and critic-based filtering, as discussed in Section 4. In total $600$ instances were annotated, equally divided into different context combinations for mind maps and table generation. Annotators consistently answered correctly across all context combinations (Appendix A), suggesting that the level of context did not significantly impact their accuracy. Figure 5 shows the overall results. The plots show 95% confidence interval of time taken by the annotators when using different modality:
@@
We note that $\langle q,s+t\rangle$ performs worse than $\langle q,s\rangle$. We believe this is due to the fact that the annotators cross-checked the answer from both the modalities, leading to increase in time to answer the question.
9 Data Generation statistics
@@
10 Limitations of Auto-QA
We propose Auto-QA as a coverage metric, in order to measure how much information from the original passage is retained in the StructSum. Note that the metric is not intended as a substitute for evaluations of overall quality. The primary benefit of Auto-QA is the synthetic question generation component, which can be run at scale without costly human annotations. In Table 4, we measured how closely the synthetic Auto-QA aligns with human generated question-answer pairs, revealing certain limitations of the metric. Notably the quality of generated questions is not always reasonable, which is mitigated using specialized filters, as discussed in Section 5. Although we generate several question-answer pairs, it is possible that none pass the filters, which we see for fewer than 10% of input passages. Such instances adversely affect the Auto-QA coverage score, contributing to the score discrepancies in Table 4. When the analysis is restricted to those subsets where the Auto-QA filtering is successful, the observed differences are diminished and fall within the bounds of experimental noise, as reported in Table 5.
11 Conclusion
In this work we study the potential of structured representations like tables and mind maps to enhance information comprehension. Utilizing our divide-and-generate prompting and iterative expansion, we achieved significant improvements in output quality (+$37pp$ for mind maps, +$15pp$ for tables) using structure-specific prompts and critics. We proposed Auto-QA based coverage metric that automatically generates QA pairs from the input text and uses StructSum outputs to answer them.
12 Acknowledgment
We would like to thank Massimo Nicosia and Yasemin Altun for their feedback and comments on the paper. We are grateful to all the volunteers who participated in our annotation experiments.
13 Limitations
We outline the limitations of our work to ensure transparency and inspire future research. First, the structured output representations we experimented with are limited to tables and mind maps. However, to comprehensively evaluate the effectiveness of our critics and pipeline, it is desirable to also evaluate other input and output modalities, e.g. image and video, considering the recent advances in VLMs. Secondly, our work and experimental findings are limited to only English sources. We plan to also explore multilingual structured summaries in future work. Third, we would to warn against the risk of blindly trusting models to generate structured summaries from an input accurately. Although we take extra care to increase the factuality of the outputs via the use of critics, and experimentally validate QA coverage, we believe that special care should be taken to verify outputs in accuracy-sensitive applications. Finally, our StructSum generation is performed using a LLM with fixed prompts, however, prior work have shown a reasonable portability of prompts across similar models (Zhou et al., 2023; Khot et al., 2023).
Despite these limitations, our work serves as an initial step in constructing reliable structured summarization evaluations, models and applications. We hope future research can greatly benefit from this starting point.
Appendix A User Study
Usually, mind maps are represented as a graph as shown in Figure 7. However, for the text comprehension user study described in Section 8.7, to avoid bias due to color or orientation, we simplify the representation as a tree (Figure 10). To establish the known query intent, annotators’ are first shown with input question, e.g., Figure 8. Next, on clicking Show content button, annotators are shown context in the form of either text (Figure 9), structure (Figure 10), or structure + text (Figure 11). The question-answer pairs were generated automatically conditioned on input text (Section 5). Annotators were also allowed to mark an instance un-answerable. The user study for tables is performed in a similar manner. We annotated $100$ question-answer pairs for both mind maps and tables. Each input instance is annotated with three different context combinations, leading to $600$ total annotations. We filtered instances that were marked un-answerable by the annotators (32% and 22% for tables and mind map study resp.). To avoid penalizing for spelling errors or other typing mistakes, the answers were evaluated via human evaluation.
We adopt timed-comprehension for answering free-form questions as a proxy measure for usefulness of generated structured representation. This makes it different from categorical data annotations. We calculate rater agreement for the questionnaire responses. We find that 89.9% of questions had full rater agreement regarding the correct response, with the only differences in the time taken to respond. Each question was shown to three raters. Table 7 shows the overall accuracy as percentage of questions answered correctly in different context. Irrespective of context combinations, annotators were able to answer the questions correctly with a high accuracy.